XluisX:

A triangle has a perimeter of 50. If 2 of its sides are equal and the third side is 5 more than the equal sides, what is the length of the third side?

XluisX:

I know im doing hw on my birthday rip

justus:

x for both sides x+5 = the third side x+x+x+5=50 3x+5=50 3x=45 x=15 x+5=20

justus:

20 is the length for the 3rd side.

TETSXPREME:

Hope You Had GREAT! Birthday @XluisX <3 no gomo

hxesmadd:

hey

hxesmadd:

happy birthday byee

Westa106:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XluisX I know im doing hw on my birthday rip \(\color{#0cbb34}{\text{End of Quote}}\) Happy bdsy